# Rotary Outliers and Rotary Offset Features in Large Language Models

### v1

@@ Abstract
Transformer-based Large Language Models (LLMs) rely on positional encodings to provide sequence position information to their attention mechanism. Rotary Positional Encodings (RoPE), which encode relative position by rotating queries and keys, have become widely used in modern LLMs. We study the features and patterns that emerge in queries and keys when using rotary embeddings. Our analysis reveals consistent patterns within the same model across layers and attention heads and across different models and architectures. We present and apply analysis techniques and show how the queries and keys use RoPE to construct various attention patterns, including attention sinks. We find and analyze outliers across models in queries and keys and find that they are likely to be found in rotary features with partial cycles. We derive bounds that tell us what rotary frequencies are likely to be selected as outlier features and at what minimum angle the query-key rotary pairs in these features tend to be above and verify the bounds empirically with models of significant architectural differences.
@@ 1 Introduction
Figure 1: Maximum query and key absolute values across all layers in Phi-1. The white dashed line separates rotary (left) from non-rotary (right) features, we further separate the first dimensions of the rotary pairs (left) from the second dimensions (right) with a red dashed line. The two bands of large absolute values are located at the pair of rotary dimensions that are rotated by the same low frequencies in a sliced rotary implementation layout. See Appendix Figure 9 for mean magnitudes.
@@
In this paper, we analyze the internal representations of models using rotary positional embeddings (RoPE) to better understand how positional information is encoded and utilized. We find a heavy concentration of large magnitudes in the queries and keys at specific rotary features (see Figure 1). Through analysis and empirical validation, we derive and verify bounds that characterize these outlier features and their effects on attention scores.
These outlier features are particularly relevant in the context of model quantization. Work by (Dettmers et al., 2022) has shown that activation outlier features can significantly impact model performance during quantization. While their analysis focused primarily on activation outliers in additive positional encodings, similar outlier features exist for networks with rotary positional embeddings. The high-norm features we identify in rotary embeddings could affect the quantization errors for activation quantized models. This is because high-norm features tend to dominate the representation space, leading to larger quantization errors when the numerical precision is reduced. Furthermore, since keys are affected by these outliers, quantized key caches may require higher precision or more complex quantization algorithms (Liu et al., 2024b) to maintain model accuracy.
Throughout the analysis we choose to ignore all non-rotary features and focus solely on the rotary features. We analyze the features and patterns that emerge in queries and keys when using rotary embeddings and make the following contributions:
  • We find outlier features appear in certain rotary frequencies across layers and heads within networks and that they exist across language models of different sizes and architectures.
  • We show how a simple upper bound on the rotary frequency can be used to predict which frequencies are likely to contain outlier activations. This upper bound depends only on context length and rotary frequency.
  • We derive a novel and simple lower bound for the query-key angles that also only depends on context length and rotary frequency. We demonstrate empirically that on average rotary outliers consistently operate within these bounds.
  • We present other analysis on rotary features, such as the mechanism by which attention sinks often emerge from the rotary outlier features identified by our bounds.
@@ 2.2 Rotary Positional Embeddings
where $\theta_{i}=c^{-2i/d}$ controls the frequency for dimension $i$, $i$ ranges from 00 to $\frac{r}{2}$ where $r$ is the number of rotary feature dimensions, and $c$ is a hyperparameter. Note that the step size $\theta_{i}$ and frequency decrease as $i$ increases. For query vector $\mathbf{q}$ and key vector $\mathbf{k}$ at position $m$, rotary embeddings transform pairs of features [^1] Note that there are several possible layouts for RoPE.:
@@
This shows that the inner product depends only on the relative position $(m-n)$ between the query and key, enabling the model to attend based on token distances. For networks with partial rotary embeddings, only a subset of the query and key feature dimensions receive rotations, with $r<d$.
@@ 3.1 Decomposing positional attention patterns
where $\phi_{i}\in(0,2\pi]$ is the initial counterclockwise angle between $\bar{\mathbf{q}}_{\langle i\rangle}$ and $\bar{\mathbf{k}}_{\langle i\rangle}$. This gives us an $\frac{r}{2}$-dimensional vector function $\mathbf{d}$ where each index $i$ contains the dot product between the counter-clockwise rotated mean query and the mean key vector at a positional distance $p$. In the rest of this paper, when we mention ”radius”, we will be speaking about the rotary pair’s mean vector’s magnitude $\|\bar{\mathbf{x}}_{\langle i\rangle}\|$.
We find that rotary feature pairs with larger mean radius tend to have more stable angles with respect to the origin (Figure 2) so we expect this simplification to preserve the key patterns as the noisier a feature’s angle is the less likely it is to contribute significantly to the dot product. However, some rotary feature vectors deviate significantly from the mean, often serving as attention sinks and this simplification will not capture these effects.
@@
From this point on, we will use ”feature” or ”rotary feature” to refer to rotary feature pairs, since we will solely speak about rotary feature pairs.
@@ 3.2 Rotary Offset Features and Bounds
What we find is that most of the outliers in rotary features are concentrated in rotary frequencies which never complete a full period. Furthermore, the ones that turn into large radii features within low frequencies tend to have a U-shaped contribution to $D(p)$, starting at a greater point than in the middle and often ending at a lesser point than they started. We call these features rotary offset features, and define them as:
@@
$x$$y$$\vec{q_{0}}$$\vec{q_{1}}$$\vec{k}$$2\phi-2\pi$$\phi$$2\pi-\phi$
Secondly, we know that the initial angle between query and key needs to be $\phi>\pi$, since $\cos\phi$ is increasing for $[\pi,2\pi]$ the rotary offset feature would be increasing above its initial value $d_{i}(0)$ for some $p\leq p_{max}$. To arrive at a tighter lower bound for the initial angle, we need to use that $cos(2\pi-\phi)=cos(\phi)$, and the angular change $\phi-(2\pi-\phi)$ gives us the angle at which the rotary offset feature returns to the same value as at zero distance (see Figure 3). This gives us the inequality $p_{max}\theta\leq 2\phi-2\pi$, where the left hand side is the total rotation at the context length and the right hand side is the maximum change in angle to return to the value at zero distance. Rearranging this, we arrive at a lower bound for the query-key angle:
@@
These bounds only apply to rotary offset features, which tend to be the cause of the majority of outliers in low frequency features. It does not apply to other features such as diagonal features which are typically high-frequency, fortunately these also tend to be of lower quantity and magnitude than rotary offset features as we will later observe (Figure 8).
@@ 4.1 Model Selection
We focus our analysis on Phi-1 (Gunasekar et al., 2023), due to its low amount of rotary features which simplifies communication of results, and to a lesser extent DeepSeek-v2-lite (DeepSeek-AI et al., 2024), which has significant architectural differences to Phi-1, including multiheaded latent attention and a single rotary key for all rotary query heads, and Llama-2-7B (Touvron et al., 2023). The partial rotary embeddings should let us separate semantic from positional features and we should be able to observe the effects of the rotary positional embeddings more clearly.
Model dh r pmax bias
Phi-1 64 32 2048 Yes
DeepSeek-v2-lite 192 64 163840 No
Llama-2-7B 128 128 4096 No
Table 1: Model architecture details showing head dimension ($h_{d}$), rotary dimension ($r$), context length ($p_{max}$), and whether the linear projections for keys and queries have bias.
Throughout the paper, we use zero-based indexing for layers, heads, and features and will completely ignore non-rotary features. For all statistics and plots involving queries and keys, we analyze the data before applying rotary transformations, unless otherwise noted. We use rotary first with a sliced rotary implementation layout (as typical by transformers implementation (Wolf et al., 2020)), we reorder DeepSeek’s activations [^2] DeepSeek-V2-Lite has rotary last interleaved layout to follow this layout as well.
We use the configured max position embeddings of models, $p_{max}$, as a proxy for trained context length. This assumes that the models were trained with this sequence length.
@@ 4.2 Single Head Analysis
We begin our analysis with a single head in Phi-1. We will later show that the patterns are in many other heads and layers, followed by an analysis showing the patterns exist across different architectures and language models.
@@
Various positional attention patterns can be constructed using the rotary features that complete several cycles during the context length, as they form a combination of sinusoidal functions of various amplitudes and frequencies. For instance, in heads with banded attention patterns, as in the head we have chosen to study, Figure 4, uses multiple low-to-mid frequency components (as seen in the decomposition in Figure 5) to build up a maximum attention score at distance zero and one and then quickly decays, forming a banded diagonal pattern.
Figure 5: Each rotary feature’s dot product component of $\mathbf{d}(p)$ for a range of positions $p$ and the total sum of the components showing the unnormalized attention score as distance increases. Feature 12 starts at about $-35$ at $p=0$. The positional attention pattern focuses on the first three tokens. The reconstructed positional attention pattern is in the Appendix, Figure 10.
@@
Some of the most consistent outlier features across heads and layers (and as we’ll see, across different models) are the features that go through exactly a full period or less of the RoPE rotations during the model’s context length. These tend to be selected to ensure the attention score does not return to the same value as at early distances. See Figure 5 for an example where feature 12 starts at a large negative value.
Rotary offset features are the ones that create the bands of large magnitudes in the query and key plots in Figure 1. In Phi-1, these bands are formed around feature 12. For the head we have selected for analysis, feature 12 is also the feature with greatest radius of the head’s rotary features. In Figure 7, we see that the feature has a large radius in both query and key and maintains consistent angles with respect to the origin (with some exceptions discussed later). The angle between the query and key vectors for feature 12 tend to form an angle $\phi_{12}$ which would result in a significant negative dot product. We also show the full rotation throughout the context length, which covers less than half a full period. This head is in no way unique and the angle between query and key are surprisingly similar across heads and layers for feature 12 when the radius is large (see Figure 8).
@@
Figure 7: Feature 12’s keys and queries are shown, with their respective means. The curved line shows the trajectory of the mean query rotated with frequency $\theta_{12}$ across the network’s context length, $p_{max}$. The query and key means maintain angles throughout that yield a consistently negative dot product. The two pink dots just above the origin are the attention sink keys, which align with the query angle. These are the only two tokens (out of over 800, bottom plot for top 10) that produce positive dot products, causing them to receive a large proportion of the attention weight.
@@
Figure 8: Mean radius of rotary key features plotted against the angle between mean query and mean key for each layer (colored) and head of Phi-1 (768 datapoints for each feature). Black dashed line for feature 11-15 are the soft lower bounds for the angle $\phi_{i}$ for each feature to serve as a rotary offset feature. Any feature without a dashed line are outside the bounds.
@@ 4.3.1 Phi-1
We observe that many of the relationships in the head we just analyzed are present across other layers and heads in Figure 8. Large radii features tend to be at certain angles $\phi_{i}$. It can be observed for Phi-1 that most of them are at low frequencies, with the exception of feature 0, which is a cyclic rotary feature. For Feature 0 the largest radii heads are located around $\phi=1rad$ which is equal to the angular step per position, $\theta_{0}=1$, so the query and key will tend to maximise their dot product at a distance of one, suggesting these are the features that select for the previous token.
@@
For feature 12, we see a similar angle as we did in our single head analysis and that it exists in many heads and across layers of the network. There exists heads with feature 11-15 with much larger radii than for feature 10 and below. For Phi-1’s feature 10, we calculate an upper bound of $p_{max}\theta_{10}=6.38$, which exceeds $2\pi$, this indicates that feature 10 and any higher frequency rotary features would be unsuitable as rotary offset features since they go through more than a full period in the context length. We observe a sharp increase in maximum radii between features 10 and 11, aligning with our predictions about which features can effectively serve as rotary offset features.
Initial angle of rotary offset features
We visualize the lower bounds in Figure 8. By inspection, the model seems to have learnt and agree with these bounds and the majority of large radii heads maintain angles above or very close to the lower bound for each feature.
Model Features %ROF Mean LB Radius Outlier Feat. UB Recall LB / LB-0.1 Recall
Phi-1 12288 31% 3.93 6 634 0.97 0.88 / 0.94
9 252 0.99 0.94 / 0.98
12 46 1.00 1.00 / 1.00
Llama-2-7b 65536 28% 4.19 6 1242 0.95 0.72 / 0.84
9 730 1.00 0.79 / 0.94
12 180 1.00 0.79 / 0.96
DeepSeek-V2-Lite 13824 28% 4.26 6 640 0.92 0.66 / 0.84
9 464 1.00 0.68 / 0.94
12 352 1.00 0.74 / 0.97
Table 2: Metrics for Phi-1 and Llama-2-7b models for keys, showing the number of rotary features, percentage of rotary features that could be selected as rotary offset features (%ROF), mean lower bound for the angles $\phi_{i}$ across all the rotary offset features, the radius for defining outlier features, the number of outlier features defined by the outlier radius, recall for the upper bound of outlier features, and the recall of outlier features for the lower bound / relaxed lower bound (by 0.1 rad or $\sim$6)) for the mean query-key angle. $\dagger$ For DeepSeek-V2-Lite, note that the rotary keys in these statistics are repeated 16 times since each rotary key is matched to 16 rotary queries.

### v2

# Rotary Offset Features in Large Language Models

@@ Abstract
Transformer-based Large Language Models (LLMs) rely on positional encodings to provide sequence position information to their attention mechanism. Rotary Positional Encodings (RoPE), which encode relative position by rotating queries and keys, have become widely used in modern LLMs. We study the features and patterns that emerge in queries and keys when using rotary embeddings and introduce the concept of rotary offset features. Our analysis reveals that these features, which frequently exhibit large activations and are often interpreted as outliers, arise consistently across layers, attention heads, and model architectures. We derive bounds predicting which rotary frequencies give rise to rotary offset features and the minimum angle between the query-key pairs for these features. We verify our predictions empirically across models of different sizes and architectures.
@@ 1 Introduction
Figure 1: Maximum query and key absolute values across all layers in Phi-1. The red dashed line separates rotary (left) from non-rotary (right) features, we further separate the first dimensions of the rotary pairs (left) from the second dimensions (right) with a white dashed line. The two bands of large absolute values are located at the pair of rotary dimensions that are rotated by the same low frequencies in a sliced rotary implementation layout. See Appendix Figure 9 for mean magnitudes.
@@
In this paper, we analyze the internal representations of models using rotary positional embeddings (RoPE) to better understand how positional information is encoded and utilized. Through empirical analysis across multiple model architectures, we discover that queries and keys exhibit consistently large magnitudes at specific pairs of rotary feature dimensions (see Figure 1). We develop bounds that characterize these pairs of large-magnitude features, which we term rotary offset features, and demonstrate how they influence attention patterns.
Feature magnitudes are particularly relevant in the context of model quantization. Work by (Dettmers et al., 2022) has shown that activation outlier features can significantly impact model performance during quantization. While their analysis focused primarily on activation outliers in models with additive positional encodings, similar features exist for networks with rotary positional embeddings. The high-norm features we identify in rotary embeddings could affect the quantization errors for activation quantized models. This is because high-norm features tend to dominate the representation space, leading to larger quantization errors when the numerical precision is reduced. Furthermore, since keys contain these features, quantized key caches may require higher precision or more complex quantization algorithms (Liu et al., 2024c) to maintain model accuracy.
In the analysis we choose to focus solely on rotary features. We analyze the features and patterns that emerge in queries and keys when using rotary embeddings and make the following contributions:
  • We find features with large magnitudes appear in pairs of feature dimensions at certain rotary frequencies, which we call rotary offset features, across layers and heads within networks and that they exist across language models of different sizes and architectures.
  • We show how a simple upper bound on the rotary frequency can be used to predict which frequencies are likely to contain rotary offset features. This upper bound depends only on context length and rotary frequency.
  • We derive a novel and simple lower bound for predicting the minimum angle the query-key pair has for the rotary offset features. This lower bound also only depends on context length and rotary frequency.
  • We present other properties and analysis of rotary features, such as a mechanism by which attention sinks are attended to in heads with rotary offset features and how context length extension affects low frequency rotary features in Llama-3 (Section C).
@@ 2.2 Rotary Positional Embeddings
where $\theta_{i}=c^{-2i/d}$ controls the frequency for dimension $i$, $i$ ranges from 0 to $\frac{r}{2}$ where $r$ is the number of rotary feature dimensions, and $c$ is a hyperparameter. Note that the step size $\theta_{i}$ and frequency decrease as $i$ increases. This means that features with higher index have lower frequencies. For query vector $\mathbf{q}$ and key vector $\mathbf{k}$ at position $m$, rotary embeddings transform pairs of features [^1] Note that there are several possible layouts for RoPE.:
@@
This shows that the inner product depends only on the relative position (mn)(m-n) between the query and key, enabling the model to attend based on token distances. For networks with partial rotary embeddings, only a subset of the query and key feature dimensions receive rotations, with $r<d$.
@@ 3.1 Decomposing positional attention patterns
where $\phi_{i}\in(0,2\pi]$ is the initial counterclockwise angle between $\bar{\mathbf{q}}_{\langle i\rangle}$ and $\bar{\mathbf{k}}_{\langle i\rangle}$. This gives us an $\frac{r}{2}$-dimensional vector function $\mathbf{d}$ where each index $i$ contains the dot product between the counter-clockwise rotated mean query and the mean key vector at a positional distance $p$. In the rest of this paper, when we mention ”radius”, we will be speaking about the rotary feature pair’s mean vector’s magnitude $\|\bar{\mathbf{x}}_{\langle i\rangle}\|$.
We find that rotary feature pairs with larger mean radius tend to have more stable angles with respect to the origin (Figure 2) so we expect this simplification to preserve the key patterns as the noisier a feature’s angle is the less likely it is to contribute significantly to the dot product. However, some rotary feature vectors deviate significantly from the mean, often serving as attention sinks, and this simplification will not capture these effects.
Figure 4: Each rotary feature’s dot product component of $\mathbf{d}(p)$ for a range of positions $p$ and the total sum of the components showing the unnormalized attention score as distance increases. Feature 12 starts at about 35-35 at $p=0$. The positional attention pattern focuses on the first three tokens. The reconstructed positional attention pattern can be seen in Figure 10.
@@
From this point on, we will use ”feature” or ”rotary feature” to refer to rotary feature pairs, since we will solely discuss rotary feature pairs.
@@ 3.2 Rotary Offset Features and Bounds
What we find is that most of the rotary features with large radius are concentrated in rotary frequencies which never complete a full period during the context length. Furthermore, the ones that turn into large-radius features within low frequencies tend to have a U-shaped contribution to $D(p)$, starting at a greater point than in the middle and often ending at a lesser point than they started. We call these features rotary offset features, and define them as:
@@
Secondly, we know that the initial angle between query and key needs to be $\phi>\pi$, since $\cos\phi$ is increasing for $\pi,2\pi]$ the rotary offset feature would be increasing above its initial value $d_{i}(0)$ for some $p\leq p_{max}$. To arrive at a tighter lower bound for the initial angle, we need to use that $cos(2\pi-\phi)=cos(\phi)$, and the angular change $\phi-(2\pi-\phi)$ gives us the angle at which the rotary offset feature returns to the same value as at zero distance (see Figure 5). This gives us the inequality $p_{max}\theta\leq 2\phi-2\pi$, where the left hand side is the total rotation at the context length and the right hand side is the maximum change in angle to return to the value at zero distance. Rearranging this, we arrive at a lower bound for the query-key angle:
@@
These bounds only apply to rotary offset features, which tend to be the cause of the majority of rotary features with large radius.
@@ 4.1 Model Selection
We focus our analysis on Phi-1 (Gunasekar et al., 2023), due to its low amount of rotary features which simplifies communication of results, and to a lesser extent DeepSeek-V2-Lite (Liu et al., 2024a), which has significant architectural differences to Phi-1, including multiheaded latent attention and a single rotary key for all rotary query heads, and Llama-3-8B/70B (Dubey et al., 2024). The partial rotary embeddings should let us separate semantic from positional features and we should be able to observe the effects of the rotary positional embeddings more clearly.
Model dh r pmax extended
Phi-1 64 32 2048 No
Llama-3-8B 128 128 4096 No
Llama-3-70B 128 128 8192 No
DeepSeek-V2-Lite 192 64 163840 Yes
Table 1: Model architecture details showing head dimension ($d_{h}$), rotary dimension ($r$), context length ($p_{max}$), and whether the linear projections for keys and queries have bias.
Throughout the paper, we use zero-based indexing for layers, heads, and features and will ignore non-rotary features. For all statistics and plots involving queries and keys, we analyze the data before applying rotary transformations, unless otherwise noted. We use rotary first with a sliced rotary implementation layout (as in most transformers models’ implementations (Wolf et al., 2020)), we reorder DeepSeek-V2-Lite’s activations [^2] DeepSeek-V2-Lite has rotary last interleaved layout. to follow this layout as well. For this layout, the rotary features’ frequencies decrease as the feature index increases.
We use the configured max position embeddings of models, $p_{max}$, as a proxy for trained context length. This assumes that the models were trained with these sequence lengths.
@@ 4.2 Single Head Analysis
We begin our analysis with a single head [^3] Layer 2 head 4 (zero-indexed) of Phi-1. in Phi-1. We will later show that the patterns are in many other heads and layers, followed by an analysis showing the patterns exist across different architectures and sizes of language models.
@@
Various positional attention patterns can be constructed using the rotary features that complete several cycles during the context length, as they form a combination of sinusoidal functions of various amplitudes and frequencies. For instance, in heads with banded attention patterns, as in the head we have chosen to study, seen in Figure 3, uses multiple low-to-mid frequency components (as seen in the decomposition in Figure 4) to build up a maximum attention score at distance zero and one and then quickly decays, forming a banded diagonal pattern.
@@
Some of the most consistent large-radius features across heads and layers (and as we’ll see, across different models) are the features that go through exactly a full period or less of the RoPE rotations during the model’s context length. These tend to be selected to ensure the attention score does not return to the same value as at early distances. See Figure 4 for an example where feature 12 starts at a large negative value.
Rotary offset features are the ones that create the bands of large magnitudes in the query and key plots in Figure 1. In Phi-1, these bands are formed around feature 12. For the head we have selected for analysis, feature 12 is also the feature with greatest radius of the head’s rotary features. In Figure 7, we see that the feature has a large radius in both query and key and maintains consistent angles with respect to the origin (with some exceptions discussed later). The angle between the query and key vectors for feature 12 tend to form an angle $\phi_{12}$ which would result in a significant negative dot product (see Figure 12 for the other 15 features). We also show the full rotation throughout the context length, which covers less than half a full period. This head is not unique and the angle between query and key are surprisingly similar for features with large radius with index 12 across different heads and layers (see Figure 8).
@@
Figure 7: Feature 12’s keys and queries are shown, with their respective means. The curved line shows the trajectory of the mean query rotated with frequency $\theta_{12}$ across the network’s context length, $p_{max}$. The query and key means maintain angles that yield a consistently negative dot product. The two pink dots just above the origin are the attention sink keys, which align with the query angle. In the bottom plot, we show that these are the only two tokens (out of over 800) that produce positive dot products, causing them to receive a large proportion of the attention weight.
@@
Figure 8: Mean radius of rotary key features plotted against the angle between mean query and mean key for each layer (colored) and head of Phi-1 (768 datapoints for each feature). Black dashed line for feature 11-15 are the lower bounds for the angle $\phi_{i}$ for each feature to serve as a rotary offset feature. Any feature without a dashed line are outside the bounds.
@@ 4.3.1 Phi-1
We observe that many of the relationships in the head we just analyzed are present across other layers and heads in Figure 8 (and across data samples 14). Large-radius features tend to be at certain angles $\phi_{i}$. It can be observed for Phi-1 that most of them are at low frequencies, with the exception of feature 0, which is a cyclic rotary feature. For Feature 0 the largest radii heads are located around $\phi=1rad$ which is equal to the angular step per position, $\theta_{0}=1$, so the query and key will tend to maximise their dot product at a distance of one, suggesting these are features that select for the previous token.
@@
For Phi-1’s feature 10, we calculate an upper bound of $p_{max}\theta_{10}=6.38$, which exceeds $2\pi$, this indicates that feature 10 and any higher frequency rotary features would be unsuitable as rotary offset features. We observe a sharp increase in maximum radii between features 10 and 11, aligning with our predictions about which features can effectively serve as rotary offset features.
Initial query-key angle of rotary offset features
We visualize the lower bounds in Figure 8. By inspection, the model seems to agree with these bounds and the majority of large-radius features maintain angles above or very close to the lower bound for each rotary frequency.
Model Features %ROF Mean LB Min. Radius Positives UB Recall LB / LB-0.1 Recall
Phi-1 12288 31% 3.93 6.0 634 0.97 0.88 / 0.94
9.0 252 0.99 0.94 / 0.98
12.0 46 1.00 1.00 / 1.00
Llama-3-8B 65536 45% 3.72 6.0 1888 0.90 0.63 / 0.78
9.0 812 1.00 0.82 / 0.94
12.0 344 1.00 0.84 / 0.99
Llama-3-70B 327680 45% 3.72 6.0 14832 0.63 0.40 / 0.52
9.0 5440 0.85 0.55 / 0.73
12.0 2024 0.97 0.63 / 0.82
DeepSeek-V2-Lite 13824 28% 4.26 6.0 640 0.92 0.66 / 0.84
9.0 464 1.00 0.68 / 0.94
12.0 352 1.00 0.74 / 0.97
Table 2: Metrics for various models for keys, showing the number of rotary features, percentage of rotary features that could be selected as rotary offset features (%ROF), mean lower bound for the angles $\phi_{i}$ across all the rotary offset features (Mean LB), the minimum radius of rotary feature pairs for defining positives (Min. Radius), the number of features considered positives (Positives) for a particular minimum radius, recall of positives for the upper bound (UB Recall), and the recall of positives for the lower bound / relaxed lower bound (by 0.1 rad or $\sim$6)) (LB / LB-0.1 Recall) for the mean query-key angle. $\dagger$ For DeepSeek-V2-Lite, note that the rotary keys in these statistics are repeated 16 times since each rotary key is matched to 16 rotary queries.